how do you solve an ellipse equation like this? when there a numbers in place of the a^2 and b^2?

(x-h)^2 (y-k)^2
_______ + _____
a^2 b^2

Answer by stanbon(48581)

(x-h)^2 (y-k)^2
_______ + _____ = 1
a^2 b^2
----------------------------------
If a^2 = 4 and b^2=9
the a = 2 and b =3
------------------------
Then the vertical axis of the ellipse has length 2*3=6
and the horizontal axis of the ellipse has length 2*2=4
---------------------------
In this case the vertical axis is the major axis
the horizontal axis is the minor axis.
---------------------------
If the numbers are switched the horizontal axis =6 is major
and the vertical axis =4 is minor.
========================================
The center in both cases is (h,k)
========================================
Cheers,
Stan H.